OpenStudy (anonymous):

Solve the inequality –2(x – 3) ≥ 5 – (x + 3) and show your work plz.

OpenStudy (anonymous):

lets start with the distributive law \[-2x+6\geq 5-x-3\]

OpenStudy (anonymous):

combine like terms on the right \[-2x+6\geq 2-x\] add \(2x\) to both sides \[6\geq 2+x\] subtract 2 from both sides \[4\geq x\] and done

OpenStudy (anonymous):

if you prefer you can write \(x\leq 4\) no difference